### v2

# Implementing Errors on Errors: Bayesian vs Frequentist

Abstract
When combining apparently inconsistent experimental results, one often implements errors on errors. The Particle Data Group’s phenomenological prescription offers a practical solution but lacks a firm theoretical foundation. To address this, D’Agostini and Cowan have proposed Bayesian and frequentist approaches, respectively, both introducing gamma-distributed auxiliary variables to model uncertainty in quoted errors. In this Letter, we show that these two formulations admit a parameter-by-parameter correspondence, and are structurally equivalent. This identification clarifies how Bayesian prior choices can be interpreted in terms of frequentist sampling assumptions, providing a unified probabilistic framework for modeling uncertainty in quoted variances.
1 Introduction
The combination of inconsistent experimental results is a recurring challenge in particle physics, particularly when managing and interpreting uncertainties. This issue arises when discrepancies between experimental data exceed the quoted statistical and systematic uncertainties, necessitating the implementation of additional uncertainty corrections.
The Particle Data Group (PDG) addresses this by introducing a phenomenological scale factor applied to the combined uncertainty:
$S:=\max\mathchoice{\!}{\!}{}{}\left(1,\,\sqrt{\chi^{2}\over N-1}\right),$ (1)
where $N$ denotes the number of experimental results being combined and $\chi^{2}$ is the chi-squared value for the combination [1]. While this method is practical and widely used, it lacks a clear theoretical foundation, leaving room for more principled statistical treatments.
In Refs. [2] and [3], D’Agostini and Cowan independently proposed approaches to address the so-called “errors on errors”—uncertainties in the quoted systematic uncertainties—based on Bayesian and frequentist inference, respectively. Both methods extend the standard profile likelihood by introducing auxiliary variables to model the uncertainty in the quoted errors, typically using gamma distributions.
While these ideas are individually known, the structural correspondence between the Bayesian and frequentist formulations has not been explicitly established in the literature. In this Letter, we show that the two approaches can in fact be mapped onto each other parameter by parameter. In particular, we demonstrate that a specific parameter choice in the Bayesian prior—often presented as an ad hoc recommendation—can be naturally understood as arising from a frequentist sampling model involving virtual repetitions of each experiment. This correspondence not only clarifies the origin of the so-called “democratic skepticism” advocated by D’Agostini, but also provides a concrete rationale for adopting symmetric gamma priors with shape and rate parameters set equal.
Our aim is not to introduce a new statistical model, but rather to clarify, connect, and unify existing approaches by revealing their structural equivalence. By identifying a parameter-by-parameter correspondence between the Bayesian and frequentist formulations, we provide a coherent probabilistic framework for modeling uncertainty in quoted variances. This perspective helps elucidate the rationale behind specific prior choices in Bayesian methods and may provide useful insight into empirical practices such as uncertainty rescaling in data combinations.
This Letter is organized as follows. In Sec. 2, we review D’Agostini’s Bayesian approach to incorporating errors on errors. In Sec. 3, we present Cowan’s frequentist formulation, recast in a form aligned with D’Agostini’s approach. In Sec. 4, we compare the two and demonstrate their structural correspondence. In Sec. 5, we summarize our findings and discuss their implications.
2 D’Agostini’s Bayesian approach
We review D’Agostini’s Bayesian approach for the errors on errors [2].
2.1 Ordinary procedure derived from Bayesian inference
Bayesian inference is based on the conditional probability111 Eq. (2) follows from $P\mathchoice{\!}{\!}{}{}\left(A\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}B\right)={P\mathchoice{\!}{\!}{}{}\left(B\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}A\right)P\mathchoice{\!}{\!}{}{}\left(A\right)\over P\mathchoice{\!}{\!}{}{}\left(B\right)}$, where $P\mathchoice{\!}{\!}{}{}\left(A\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}B\right):={P\mathchoice{\!}{\!}{}{}\left(A\cap B\right)\over P\mathchoice{\!}{\!}{}{}\left(B\right)}$.
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}}\right)$ $\propto f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu}\right)f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right),$ (2)
where
${\color[rgb]{0,0.5,1}\boldsymbol{d}}:=\Set{{\color[rgb]{0,0.5,1}d_{i}}}_{i=1,\dots,N}=\Set{{\color[rgb]{0,0.5,1}d_{1}},\dots,{\color[rgb]{0,0.5,1}d_{N}}}$ (3)
is the data set, with roman indices $i,j,\dots$ running from $1$ to $N$ to label the experimental data, typically obtained from different experimental collaborations. Here, $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu}\right)$ represents the likelihood of observing $\color[rgb]{0,0.5,1}\boldsymbol{d}$} under the hypothesis that the true value is $\color[rgb]{1,0.2,0}\mu$}, and $f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)$ is the prior probability density function (PDF) of $\color[rgb]{1,0.2,0}\mu$}. In Eq. (2), both $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}}\right)$ and $f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)$ describe theoretical distributions for the “true” value $\color[rgb]{1,0.2,0}\mu$}, while $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu}\right)$ represents the experimental probability of obtaining the dataset $\color[rgb]{0,0.5,1}\boldsymbol{d}$}, assuming a true value $\color[rgb]{1,0.2,0}\mu$}.
Random variables Fixed quantities
Theoretical $^{\color[rgb]{1,0.2,0}\hat{\mu}}$, $^{\color[rgb]{1,0.2,0}\boldsymbol{\hat{\sigma}}}$, $^{\color[rgb]{1,0.2,0}\boldsymbol{\hat{r}}}$, $^{\color[rgb]{1,0.2,0}\boldsymbol{\hat{\omega}}}$ $\color[rgb]{1,0.2,0}\mu$}, $\color[rgb]{1,0.2,0}\boldsymbol{\sigma}$}, $\color[rgb]{1,0.2,0}\boldsymbol{r}$}, $\color[rgb]{1,0.2,0}\boldsymbol{\omega}$}, $\color[rgb]{1,0.2,0}\boldsymbol{v}$}
Experimental $^{\color[rgb]{0,0.5,1}\boldsymbol{\hat{d}}}$, $^{\color[rgb]{0,0.5,1}\boldsymbol{\hat{w}}}$, $^{\color[rgb]{0,0.5,1}\boldsymbol{\hat{s}}}$ $\color[rgb]{0,0.5,1}\boldsymbol{d}$}, $\color[rgb]{0,0.5,1}\boldsymbol{w}$}, $\color[rgb]{0,0.5,1}\boldsymbol{s}$}
Table 1: Variables used in this work, arranged by whether they are theoretical or experimental (rows) and whether they are random variables or fixed quantities (columns). Nuisance parameters are $\color[rgb]{0.03,1,0.25}k$}, $\color[rgb]{0.03,1,0.25}\lambda$}, $\color[rgb]{0.03,1,0.25}n$}, and $\color[rgb]{0.03,1,0.25}\varepsilon$}.
In the usual frequentist approach, each experimental result $\color[rgb]{0,0.5,1}\hat{d}_{i}$} in the set ${\color[rgb]{0,0.5,1}\boldsymbol{\hat{d}}}:=\set{{\color[rgb]{0,0.5,1}\hat{d}_{1}},\dots,{\color[rgb]{0,0.5,1}\hat{d}_{N}}}$ is treated as a random variable independently distributed around the fixed true value $\color[rgb]{1,0.2,0}\mu$}. Here and hereafter, a hat $\hat{}\,$ denotes all estimators and other random variables derived from data, regardless of whether they appear in a frequentist or Bayesian framework. This notation distinguishes the hatted random variables from fixed parameters such as $\color[rgb]{1,0.2,0}\mu$} and from deterministic values such as $\color[rgb]{0,0.5,1}\boldsymbol{d}$}. These distinctions are summarized in Table 1.
Each $\color[rgb]{0,0.5,1}\hat{d}_{i}$} is typically modeled as following a normal distribution with mean $\color[rgb]{1,0.2,0}\mu$} and standard deviation $\color[rgb]{1,0.2,0}\sigma_{i}$}:
$\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{1,0.2,0}\sigma_{i}^{2}}\right):={1\over\sqrt{2\pi}{\color[rgb]{1,0.2,0}\sigma_{i}}}\exp\mathchoice{\!}{\!}{}{}\left[-{\left({\color[rgb]{0,0.5,1}d_{i}}-{\color[rgb]{1,0.2,0}\mu}\right)^{2}\over 2{\color[rgb]{1,0.2,0}\sigma_{i}^{2}}}\right].$ (4)
We also collectively write ${\color[rgb]{1,0.2,0}\boldsymbol{\sigma}}:=\Set{{\color[rgb]{1,0.2,0}\sigma_{i}}}_{i=1,\dots,N}$. Under the above assumption, the PDF for $^{\color[rgb]{0,0.5,1}\boldsymbol{\hat{d}}}$ is given by
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu}\right)=f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{1,0.2,0}\boldsymbol{\sigma}}\right):=\prod_{i}\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{1,0.2,0}\sigma_{i}^{2}}\right).$ (5)
Here and hereafter, we use the shorthand notation $\prod_{i}:=\prod_{i=1}^{N}$, $\sum_{i}:=\sum_{i=1}^{N}$, etc.
Now we come back to the Bayesian approach. When one further assumes that the prior is uniform, $f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)=\text{const.}$,222 The uniform prior adopted here is improper and non-integrable, but this does not affect our subsequent analysis as long as the likelihood function (7) decays sufficiently fast. and that $\color[rgb]{1,0.2,0}\boldsymbol{\sigma}$} are exactly equal to the quoted stated uncertainties by the experimental collaborations,
${\color[rgb]{0,0.5,1}\boldsymbol{s}}:=\Set{{\color[rgb]{0,0.5,1}s_{1}},\dots,{\color[rgb]{0,0.5,1}s_{N}}}$, (6)
one ends up with the ordinary procedure:
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{1,0.2,0}\boldsymbol{\sigma}},f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)=\text{const.}\right)$ $\propto\prod_{i}\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{0,0.5,1}s_{i}^{2}}\right)\propto\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\overline{d}},{\color[rgb]{0,0.5,1}\overline{s}^{2}}\right),$ (7)
where
${\color[rgb]{0,0.5,1}\overline{s}}$ $:={1\over\sqrt{\sum_{i}{1\over{\color[rgb]{0,0.5,1}s_{i}^{2}}}}},$ ${\color[rgb]{0,0.5,1}\overline{d}}$ $:={\sum_{i}{{\color[rgb]{0,0.5,1}d_{i}}\over{\color[rgb]{0,0.5,1}s_{i}^{2}}}\over\sum_{j}{1\over{\color[rgb]{0,0.5,1}s_{j}^{2}}}}={\color[rgb]{0,0.5,1}\overline{s}^{2}}\sum_{i}{{\color[rgb]{0,0.5,1}d_{i}}\over{\color[rgb]{0,0.5,1}s_{i}^{2}}}.$ (8)
Throughout this Letter, an overbar denotes an averaged quantity. Accordingly, the expected value, the variance, and the standard deviation are
$\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\mu}}\right$] $:=\int\text{d}{\color[rgb]{1,0.2,0}\mu}\,{\color[rgb]{1,0.2,0}\mu}\,\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\overline{d}},{\color[rgb]{0,0.5,1}\overline{s}^{2}}\right)={\color[rgb]{0,0.5,1}\overline{d}},$ (9)
$\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\mu}}\right$] $:=\operatorname{E}\mathchoice{\!}{\!}{}{}\left[\left({\color[rgb]{1,0.2,0}\hat{\mu}}-\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\mu}}\right]\right)^{2}\right]=\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\mu}^{2}}\right]-\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\mu}}\right]={\color[rgb]{0,0.5,1}\overline{s}^{2}},$ (10)
$\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\mu}}\right$] $:=\sqrt{\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\mu}}\right]}={\color[rgb]{0,0.5,1}\overline{s}},$ (11)
for the normal distribution.
2.2 Probabilistic modeling of skepticism
In the above, we have first treated the standard deviations $\color[rgb]{1,0.2,0}\boldsymbol{\sigma}$} as unknown quantities, and then assumed them to be fixed at $\color[rgb]{0,0.5,1}\boldsymbol{s}$}, namely, ${\color[rgb]{1,0.2,0}\sigma_{i}}={\color[rgb]{0,0.5,1}s_{i}}$. Therefore, it is natural to promote $\color[rgb]{1,0.2,0}\boldsymbol{\sigma}$} to a set of random variables:
${\color[rgb]{1,0.2,0}\boldsymbol{\hat{\sigma}}}=\Set{{\color[rgb]{1,0.2,0}\hat{\sigma}_{1}},\dots,{\color[rgb]{1,0.2,0}\hat{\sigma}_{N}}}$. (12)
D’Agostini introduces the following ratios as alternative random variables to $^{\color[rgb]{1,0.2,0}\boldsymbol{\hat{\sigma}}}$:
${\color[rgb]{1,0.2,0}\hat{r}_{i}}$ $:={{\color[rgb]{1,0.2,0}\hat{\sigma}_{i}}\over{\color[rgb]{0,0.5,1}s_{i}}}.$ (13)
The basic assumption is that the inverse squares of these ratios
${\color[rgb]{1,0.2,0}\hat{\omega}_{i}}$ $:={1\over{\color[rgb]{1,0.2,0}\hat{r}_{i}^{2}}}={{\color[rgb]{0,0.5,1}s_{i}^{2}}\over{\color[rgb]{1,0.2,0}\hat{\sigma}_{i}^{2}}}$ (14)
follow the gamma distribution, whose PDF is given by333 We follow the standard convention in statistics, where the shape parameter is listed first, followed by the rate parameter. This contrasts with the convention used by the PDG [1], where the rate parameter appears first.
$g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\omega_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ $:={{\color[rgb]{0.03,1,0.25}\lambda^{k}}\over\Gamma\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k}\right)}{\color[rgb]{1,0.2,0}\omega_{i}^{{\color[rgb]{0.03,1,0.25}k}-1}}e^{-{\color[rgb]{0.03,1,0.25}\lambda}{\color[rgb]{1,0.2,0}\omega_{i}}},$ (15)
where ${\color[rgb]{0.03,1,0.25}k}>0$ and ${\color[rgb]{0.03,1,0.25}\lambda}>0$ are shape and rate parameters, respectively, which are treated as nuisance parameters. [^4] D’Agostini first introduces $i$-dependent $\color[rgb]{0.03,1,0.25}k_{i}$} and $\color[rgb]{0.03,1,0.25}\lambda_{i}$}, and later assumes that $\color[rgb]{1,0.2,0}\hat{\omega}_{i}$} are identically distributed for all $i$, namely, ${\color[rgb]{0.03,1,0.25}k_{i}}={\color[rgb]{0.03,1,0.25}k}$ and ${\color[rgb]{0.03,1,0.25}\lambda_{i}}={\color[rgb]{0.03,1,0.25}\lambda}$, calling it democratic skepticism. We adopt this assumption from the beginning, since the $i$-dependence can be trivially recovered if needed.
The gamma distribution satisfies the following scaling property:
$\alpha g\mathchoice{\!}{\!}{}{}\left(\alpha{\color[rgb]{1,0.2,0}\omega_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ $=g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\omega_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},\alpha{\color[rgb]{0.03,1,0.25}\lambda}\right),$ (16)
with $\alpha>0$. For later convenience, we also list its expected value and variance:
$\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]$ $=\int_{0}^{\infty}\text{d}{\color[rgb]{1,0.2,0}\omega_{i}}\,{\color[rgb]{1,0.2,0}\omega_{i}}\,g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\omega_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)={{\color[rgb]{0.03,1,0.25}k}\over{\color[rgb]{0.03,1,0.25}\lambda}},$ (17)
$\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]$ $=\operatorname{E}\mathchoice{\!}{\!}{}{}\left[\left({\color[rgb]{1,0.2,0}\hat{\omega}_{i}}-\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]\right)^{2}\right]=\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}^{2}}\right]-\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]={{\color[rgb]{0.03,1,0.25}k}\over{\color[rgb]{0.03,1,0.25}\lambda^{2}}}.$ (18)
They both exist for all admissible ${\color[rgb]{0.03,1,0.25}\lambda}>0$.
One of the main purposes of this Letter is to provide a clearer theoretical foundation for the somewhat arbitrary assumption (15) by examining how it aligns with Cowan’s frequentist approach.
2.3 D’Agostini’s ratio
Finally, taking into account the Jacobian in the probability element, we obtain the PDF in terms of the ratio:
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}r_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)\text{d}{\color[rgb]{1,0.2,0}r_{i}}$ $=g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\omega_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)\text{d}{\color[rgb]{1,0.2,0}\omega_{i}}={2\over{\color[rgb]{1,0.2,0}r_{i}^{3}}}\,g\mathchoice{\!}{\!}{}{}\left({1\over{\color[rgb]{1,0.2,0}r_{i}^{2}}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)\text{d}{\color[rgb]{1,0.2,0}r_{i}}={2{\color[rgb]{0.03,1,0.25}\lambda^{k}}\over\Gamma\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k}\right)}{e^{-{\color[rgb]{0.03,1,0.25}\lambda}/{\color[rgb]{1,0.2,0}r_{i}^{2}}}\over{\color[rgb]{1,0.2,0}r_{i}^{2{\color[rgb]{0.03,1,0.25}k}+1}}}\text{d}{\color[rgb]{1,0.2,0}r_{i}}.$ (19)
Consequently,
$\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{r}_{i}}\right]$ $=\int_{0}^{\infty}\text{d}{\color[rgb]{1,0.2,0}r_{i}}\,{\color[rgb]{1,0.2,0}r_{i}}\,f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}r_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)={{\color[rgb]{0.03,1,0.25}\sqrt{\lambda}}\,\Gamma\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k}-{1\over 2}\right)\over\Gamma\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k}\right)},$ (20)
$\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{r}_{i}}\right]$ $={\color[rgb]{0.03,1,0.25}\lambda}\left({1\over{\color[rgb]{0.03,1,0.25}k}-1}-{\Gamma^{2}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k}-{1\over 2}\right)\over\Gamma^{2}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k}\right)}\right).$ (21)
They both exist when ${\color[rgb]{0.03,1,0.25}\lambda}>0$ and ${\color[rgb]{0.03,1,0.25}k}>1$.
Figure 1: PDFs $g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\omega}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ and $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}r}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ are shown in the left and right panels, respectively. For simplicity, the $i$-dependence of $\color[rgb]{1,0.2,0}\omega_{i}$} and $\color[rgb]{1,0.2,0}r_{i}$} is omitted in this figure and caption. The nuisance parameters are set to D’Agostini’s recommended values ${\color[rgb]{0.03,1,0.25}k}=1.29$, ${\color[rgb]{0.03,1,0.25}\lambda}=0.589$ (solid), and to Cowan’s ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=1/4{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}$ with ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.2$ (dashed), $0.5$ (dotted), and 11 (dot-dashed), corresponding to ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=6.25$, $1$, and $0.25$, respectively. The associated numbers of virtual repetitions are ${\color[rgb]{0.03,1,0.25}n}=13.5$, $3$, and $1.5$. Among these, the choice ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.5$ corresponds to our version of D’Agostini’s criterion $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}}\right]=\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}}\right$].
The above expressions (20) and (21) can be used to choose natural values of $\color[rgb]{0.03,1,0.25}\lambda$} and $\color[rgb]{0.03,1,0.25}k$}. Reasonable constraints might be
$\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{r}_{i}}\right]$ $=1,$ $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{r}_{i}}\right]$ $\simeq\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{r}_{i}}\right]=1.$ (22)
The former condition follows from the assumption that the standard deviations on average agree with the stated uncertainties, while the latter provides a conservative estimate.555 As discussed in Section 2.3.6 of Ref. [4], some published measurements may contain biases or underestimate uncertainties, potentially affecting D’Agostini’s assumption that quoted uncertainties are unbiased and conservatively estimated. The constraints (22) are satisfied by
${\color[rgb]{0.03,1,0.25}k}$ $\simeq 1.29,$ ${\color[rgb]{0.03,1,0.25}\lambda}$ $\simeq 0.589.$ (23)
These are the recommended values by D’Agostini; see Fig. 1 for the corresponding PDFs.
We see that the former condition in Eq. (22) is physically motivated, reflecting the assumption that the quoted errors are unbiased, whereas the latter serves as a conservative estimate with no clear theoretical preference. As an alternative to $\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{r}_{i}}\right]=1$, one may consider imposing the condition
$\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]=1.$ (24)
Together with Eq. (17), this leads to
${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda},$ (25)
with the variance
$\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]=\frac{1}{{\color[rgb]{0.03,1,0.25}k}}.$ (26)
We will show in Section 4 that this choice arises naturally in Cowan’s frequentist formulation, offering a compelling theoretical perspective.
Further imposing the conservative estimate $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]=1$ yields
${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=1,$ (27)
for which D’Agostini’s ratio satisfies $\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{r}_{i}}\right]=\sqrt{\pi}\simeq 1.77$, while its variance diverges as $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{r}_{i}}\right]\to 1/({\color[rgb]{0.03,1,0.25}k}-1)\to\infty$.
2.4 Likelihood function
For the likelihood function $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu}\right)$ in the Bayesian inference (2), we assume that the results are independent from each other and further that $\color[rgb]{0,0.5,1}\hat{d}_{i}$} follows a normal distribution with standard deviation ${\color[rgb]{1,0.2,0}r_{i}}{\color[rgb]{0,0.5,1}s_{i}}$ ($\color[rgb]{1,0.2,0}\sigma_{i}$} in the previous notation):
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu}\right)=\prod_{i}f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{0,0.5,1}s_{i}},{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right),$ (28)
where
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{0,0.5,1}s_{i}},{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ $=\int_{0}^{\infty}\text{d}{\color[rgb]{1,0.2,0}r_{i}}\,\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},\left({\color[rgb]{1,0.2,0}r_{i}}{\color[rgb]{0,0.5,1}s_{i}}\right)^{2}\right)f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}r_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right);$ (29)
cf. Eq. (5). Putting Eqs. (4) and (19), this assumption leads to
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{0,0.5,1}s_{i}},{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ $={{\color[rgb]{0.03,1,0.25}\lambda^{k}}\over\sqrt{2\pi}{\color[rgb]{0,0.5,1}s_{i}}}{\Gamma\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k}+{1\over 2}\right)\over\Gamma\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k}\right)}\left({\color[rgb]{0.03,1,0.25}\lambda}+{\left({\color[rgb]{0,0.5,1}d_{i}}-{\color[rgb]{1,0.2,0}\mu}\right)^{2}\over 2{\color[rgb]{0,0.5,1}s_{i}^{2}}}\right)^{-\left({\color[rgb]{0.03,1,0.25}k}+{1\over 2}\right)}.$ (30)
This PDF exhibits the kernel of a Student’s $t$ distribution, reflecting marginalization over a gamma-distributed variance.666 Integrating the normal likelihood in Eq. (29) over the gamma prior yields a Student’s $t$ distribution with $\nu=2{\color[rgb]{0.03,1,0.25}k}$. Equation (30) matches its kernel up to reparameterization, exposing the heavy tails from variance uncertainty.
When one further assumes a uniform prior distribution for $^{\color[rgb]{1,0.2,0}\hat{\mu}}$, the resulting posterior takes the form
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{0,0.5,1}\boldsymbol{s}},{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)\propto f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}\boldsymbol{d}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{0,0.5,1}\boldsymbol{s}},{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)\propto\prod_{i}\left({\color[rgb]{0.03,1,0.25}\lambda}+{\left({\color[rgb]{0,0.5,1}d_{i}}-{\color[rgb]{1,0.2,0}\mu}\right)^{2}\over 2{\color[rgb]{0,0.5,1}s_{i}^{2}}}\right)^{-\left({\color[rgb]{0.03,1,0.25}k}+{1\over 2}\right)}.$ (31)
Finally, integrating out the nuisance parameters $\color[rgb]{0.03,1,0.25}k$} and $\color[rgb]{0.03,1,0.25}\lambda$}, we obtain
$f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{0,0.5,1}\boldsymbol{s}}\right)$ $=\int\text{d}{\color[rgb]{0.03,1,0.25}k}\,\text{d}{\color[rgb]{0.03,1,0.25}\lambda}\,f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{0,0.5,1}\boldsymbol{s}},{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right),$ (32)
where $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ quantifies the confidence in each possible set of parameters. Alternatively, Eq. (31) can be evaluated simply by substituting D’Agostini’s recommended values (23) or our own (25), which is arguably the only feasible approach in practice.
3 Cowan’s frequentist approach
We review Cowan’s frequentist approach to dealing with errors on errors [3].
3.1 Cowan’s sampling scheme
In the Bayesian approach, the “true value” $^{\color[rgb]{1,0.2,0}\hat{\mu}}$ is treated as a theoretical random variable that follows a prior distribution $f_{\circ}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\right)$ and is updated to a posterior distribution $f\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\mu}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0,0.5,1}\boldsymbol{d}},{\color[rgb]{0,0.5,1}\boldsymbol{s}}\right)$ upon observing data. In contrast, the frequentist approach considers $\color[rgb]{1,0.2,0}\mu$} to be the fixed “true value”, while the observed variables fluctuate around it according to Eq. (5).
To handle the “errors on errors” within the frequentist framework, Cowan provides a way to treat each estimated variance $\color[rgb]{0,0.5,1}\hat{w}_{i}$} as an experimental random variable fluctuating around a so-to-say “true variance” $\color[rgb]{1,0.2,0}v_{i}$}. To implement this idea, Cowan introduces $\color[rgb]{0.03,1,0.25}n$} (virtual) repetitions for each real experiment $i$, resulting in $\color[rgb]{0.03,1,0.25}n$} random variables for each: [^7] In principle, $\color[rgb]{0.03,1,0.25}n$} may vary across experiments, resulting in an $i$-dependent $\color[rgb]{0.03,1,0.25}n_{i}$}. However, this dependence is typically neglected in practice [5, 6], corresponding to D’Agostini’s notion of democratic skepticism; see footnote 4. Here we also omit the $i$-dependence, as it can be trivially recovered if needed.
$\Set{{\color[rgb]{0,0.5,1}\hat{d}_{i\,\textsf{j}}}}_{\textsf{j}=1,\dots,{\color[rgb]{0.03,1,0.25}n}}.$ (33)
These are assumed to follow normal distributions:
$\mathcal{N}\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}d_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}\mu},{\color[rgb]{1,0.2,0}v_{i}}\right).$ (34)
Here and hereafter, the Roman typewriter-style index j runs from 1 to $\color[rgb]{0.03,1,0.25}n$}, indexing the virtual repetitions associated with each real experiment $i$. The sample mean and the unbiased sample variance of these outcomes are defined as usual:
${\color[rgb]{0,0.5,1}\overline{\hat{d}_{i}}}$ $:={1\over{\color[rgb]{0.03,1,0.25}n}}\sum_{\textsf{j}=1}^{{\color[rgb]{0.03,1,0.25}n}}{\color[rgb]{0,0.5,1}\hat{d}_{i\,\textsf{j}}},$ ${\color[rgb]{0,0.5,1}\hat{w}_{i}}$ $:={1\over{\color[rgb]{0.03,1,0.25}n}-1}\sum_{\textsf{j}=1}^{{\color[rgb]{0.03,1,0.25}n}}\left({\color[rgb]{0,0.5,1}\hat{d}_{i\,\textsf{j}}}-{\color[rgb]{0,0.5,1}\overline{\hat{d}_{i}}}\right)^{2}.$ (35)
Given the normal distributions (34), the random variable ${{\color[rgb]{0.03,1,0.25}n}-1\over{\color[rgb]{1,0.2,0}v_{i}}}{\color[rgb]{0,0.5,1}\hat{w}_{i}}$ follows a chi-squared distribution with ${\color[rgb]{0.03,1,0.25}n}-1$ degrees of freedom, which then can be written in terms of the gamma distribution (15) as:
$\chi^{2}_{{\color[rgb]{0.03,1,0.25}n}-1}\mathchoice{\!}{\!}{}{}\left({{\color[rgb]{0.03,1,0.25}n}-1\over{\color[rgb]{1,0.2,0}v_{i}}}{\color[rgb]{0,0.5,1}w_{i}}\right)$ $=g\mathchoice{\!}{\!}{}{}\left({{\color[rgb]{0.03,1,0.25}n}-1\over{\color[rgb]{1,0.2,0}v_{i}}}{\color[rgb]{0,0.5,1}w_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{{\color[rgb]{0.03,1,0.25}n}-1\over 2},{1\over 2}\right).$ (36)
That is, the PDF for the variance is obtained from the probability elements as
$p\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}w_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}v_{i}},{\color[rgb]{0.03,1,0.25}n}\right)\text{d}{\color[rgb]{0,0.5,1}w_{i}}$ $=g\mathchoice{\!}{\!}{}{}\left({{\color[rgb]{0.03,1,0.25}n}-1\over{\color[rgb]{1,0.2,0}v_{i}}}{\color[rgb]{0,0.5,1}w_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{{\color[rgb]{0.03,1,0.25}n}-1\over 2},{1\over 2}\right)\text{d}\mathchoice{\!}{\!}{}{}\left({{\color[rgb]{0.03,1,0.25}n}-1\over{\color[rgb]{1,0.2,0}v_{i}}}{\color[rgb]{0,0.5,1}w_{i}}\right)$
$=g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}w_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{{\color[rgb]{0.03,1,0.25}n}-1\over 2},{{\color[rgb]{0.03,1,0.25}n}-1\over 2{\color[rgb]{1,0.2,0}v_{i}}}\right)\text{d}{\color[rgb]{0,0.5,1}w_{i}},$ (37)
where we used the scaling property (16). It follows that
$\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]$ $={{\color[rgb]{1,0.2,0}v_{i}}},$ $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]$ $={2\over{\color[rgb]{0.03,1,0.25}n}-1}{\color[rgb]{1,0.2,0}v_{i}^{2}},$ $\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]$ $=\sqrt{2\over{\color[rgb]{0.03,1,0.25}n}-1}\,{\color[rgb]{1,0.2,0}v_{i}}.$ (38)
3.2 Errors-on-errors parameter
Cowan defines what is called the errors-on-errors parameter:888 We follow the notation $\color[rgb]{0.03,1,0.25}\varepsilon$} used in Ref. [7], rather than $\color[rgb]{0.03,1,0.25}r$} in Ref. [3], for the same quantity (not to be confused with D’Agostini’s $\color[rgb]{1,0.2,0}r_{i}$}). Although this parameter may in principle depend on $i$, we omit the $i$-dependence while retaining the original name, “errors-on-errors,” which reflects its possible variability across experiments; see also footnotes 4 and 7.
${\color[rgb]{0.03,1,0.25}\varepsilon}$ $:={1\over 2}{\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\over\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]}={1\over\sqrt{2\left({\color[rgb]{0.03,1,0.25}n}-1\right)}}.$ (39)
The spirit behind this definition can be explained as follows [3]: For ${\color[rgb]{0,0.5,1}\hat{s}_{i}^{2}}:={\color[rgb]{0,0.5,1}\hat{w}_{i}}$, the standard error propagation gives $\Delta{\color[rgb]{0,0.5,1}w_{i}}\simeq 2{\color[rgb]{0,0.5,1}s_{i}}\Delta{\color[rgb]{0,0.5,1}s_{i}}$. Cowan then expects the analogous relation between the standard deviations: $\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\simeq 2\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]$. If one were to approximate $\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\sim\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]$, then one would vaguely expect999 Note, however, that this naive treatment leads to $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\sim 0$, implying $\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\sim 0$. This limitation can also be seen more explicitly: It suggests that ${1\over 2}{\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]\over\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]}=\frac{1}{2}\sqrt{\frac{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}^{4}}}}{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}^{2}}}^{2}}-1}$ should be close to ${\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\over\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]}=\sqrt{\frac{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}^{2}}}}{\Braket{{\color[rgb]{0,0.5,1}\hat{s}_{i}}}^{2}}-1},$ which is unjustifiable.
${\color[rgb]{0.03,1,0.25}\varepsilon}\sim{\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]\over\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{s}_{i}}\right]}.$ (40)
In this sense, $\color[rgb]{0.03,1,0.25}\varepsilon$} is said to be a measure of relative error on error.
Regardless of how it was initially motivated, the errors-on-errors parameter (39) serves as a convenient tool to express the auxiliary quantity $\color[rgb]{0.03,1,0.25}n$} in terms of an intuitive measure of relative uncertainty:
${\color[rgb]{0.03,1,0.25}n}$ $=1+\frac{1}{2{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}}.$ (41)
The resultant PDF reads
$p\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}w_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{1,0.2,0}v_{i}},1+{1\over 2{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}}\right)$ $=g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{0,0.5,1}w_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{1\over 4{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}},{1\over 4{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}{\color[rgb]{1,0.2,0}v_{i}}}\right),$ (42)
and the variance and standard deviation of $\color[rgb]{0,0.5,1}\hat{w}_{i}$} become
$\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]$ $=4{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}{\color[rgb]{1,0.2,0}v_{i}^{2}},$ $\sigma\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]=\sqrt{\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{0,0.5,1}\hat{w}_{i}}\right]}$ $=2{\color[rgb]{0.03,1,0.25}\varepsilon}{\color[rgb]{1,0.2,0}v_{i}}.$ (43)
4 Comparison between D’Agostini and Cowan
This section presents the main findings of the Letter. We compare D’Agostini’s Bayesian and Cowan’s frequentist approaches to uncertainty in quoted variances, focusing on the structure of their respective probabilistic formulations.
D’Agostini’s Bayesian approach treats $\color[rgb]{0,0.5,1}s_{i}^{2}$} and $\color[rgb]{1,0.2,0}\hat{\sigma}_{i}^{2}$} as the experimental deterministic value and the theoretical random variable, respectively, whereas Cowan’s frequentist approach treats $\color[rgb]{0,0.5,1}\hat{w}_{i}$} and $\color[rgb]{1,0.2,0}v_{i}$} as the experimental random variable and the deterministic theoretical value.
Here we point out, on physical grounds, that the ratio ${\color[rgb]{0,0.5,1}s_{i}^{2}}/{\color[rgb]{1,0.2,0}\hat{\sigma}_{i}^{2}}$ corresponds to the ratio ${\color[rgb]{0,0.5,1}\hat{w}_{i}}/{\color[rgb]{1,0.2,0}v_{i}}$:
$\underset{\text{D'Agostini's Bayesian}}{{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}={1\over{\color[rgb]{1,0.2,0}\hat{r}_{i}^{2}}}={{\color[rgb]{0,0.5,1}s_{i}^{2}}\over{\color[rgb]{1,0.2,0}\hat{\sigma}_{i}^{2}}}}$ $\stackrel{{\scriptstyle\text{correspond}}}{{\longleftrightarrow}}\underset{\text{Cowan's frequentist}}{{\color[rgb]{0,0.5,1}\hat{w}_{i}}\over{\color[rgb]{1,0.2,0}v_{i}}}.$ (44)
In Cowan’s Eq. (36) or (37), the probability element reads
$\chi^{2}_{{\color[rgb]{0.03,1,0.25}n}-1}\mathchoice{\!}{\!}{}{}\left({{\color[rgb]{0.03,1,0.25}n}-1\over{\color[rgb]{1,0.2,0}v_{i}}}{\color[rgb]{0,0.5,1}w_{i}}\right)\text{d}\mathchoice{\!}{\!}{}{}\left({{\color[rgb]{0.03,1,0.25}n}-1\over{\color[rgb]{1,0.2,0}v_{i}}}{\color[rgb]{0,0.5,1}w_{i}}\right)$ $=g\mathchoice{\!}{\!}{}{}\left({{\color[rgb]{0,0.5,1}w_{i}}\over{\color[rgb]{1,0.2,0}v_{i}}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{{\color[rgb]{0.03,1,0.25}n}-1\over 2},{{\color[rgb]{0.03,1,0.25}n}-1\over 2}\right)\text{d}{{\color[rgb]{0,0.5,1}w_{i}}\over{\color[rgb]{1,0.2,0}v_{i}}},$ (45)
where we have again used the scaling property (16). This can be directly compared with D’Agostini’s gamma distribution in Eq. (15):
$g\mathchoice{\!}{\!}{}{}\left({\color[rgb]{1,0.2,0}\omega_{i}}\mathchoice{\;}{\;}{\,}{\,}\middle|\mathchoice{\;}{\;}{\,}{\,}{\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)\text{d}{\color[rgb]{1,0.2,0}\omega_{i}}.$ (46)
The shape and rate parameters can thus be identified as
${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}={{\color[rgb]{0.03,1,0.25}n}-1\over 2}={1\over 4{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}}.$ (47)
We see that the choice ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}$, as suggested in Eq. (25), provides a better match with the result from Cowan’s frequentist approach.101010 Cowan also touches on related Bayesian approaches in his talk [8], noting that choosing ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}$ yields a gamma distribution with unit expectation and adjustable variance. His reference to the inverse gamma distribution, in comparison with D’Agostini’s approach, suggests a perspective not fully aligned with the formulation considered here, which relies exclusively on the gamma distribution. The corresponding PDFs are shown in Fig. 1 with the values ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.2$, $0.5$, and $1$, namely, ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=6.25$, $1$, and $0.25$, respectively. Accordingly, the corresponding numbers of virtual repetitions are ${\color[rgb]{0.03,1,0.25}n}=13.5$, $3$, and $1.5$, respectively. Among these, the choice ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.5$ corresponds to our version of D’Agostini’s criterion $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]=\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]$, leading to Eq. (27).
5 Summary and discussion
In this Letter, we have clarified the structural correspondence between D’Agostini’s Bayesian and Cowan’s frequentist approaches to implementing errors on errors. Both methods enhance the conventional treatment of experimental uncertainties by modeling auxiliary random variables with gamma distributions. Although their starting points differ—Bayesian priors versus frequentist sampling models—we have established a parameter-by-parameter mapping between them.
In particular, we have demonstrated that the Bayesian assumption of a symmetric gamma prior with shape and rate parameters $\left({\color[rgb]{0.03,1,0.25}k},{\color[rgb]{0.03,1,0.25}\lambda}\right)$ corresponds naturally to Cowan’s frequentist model with $\color[rgb]{0.03,1,0.25}n$} virtual repetitions, where ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=\left({\color[rgb]{0.03,1,0.25}n}-1\right)/2=1/4{\color[rgb]{0.03,1,0.25}\varepsilon^{2}}$. This identification provides a theoretical rationale for the prior choice ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}$, ensuring $\operatorname{E}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]=1$ and $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]=1/{\color[rgb]{0.03,1,0.25}k}$, and provides a basis for choosing specific values such as ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=1$, which satisfy $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]=\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]$ as our version of D’Agostini’s recommendation.
Our aim has not been to propose a new statistical model, but rather to unify existing approaches by revealing their underlying equivalence. This correspondence offers a principled perspective that sheds light on empirical procedures such as uncertainty rescaling, and clarifies how specific prior choices in Bayesian methods can be interpreted in terms of frequentist assumptions about sampling variability.
We hope that the parameter-by-parameter identification presented here serves to bridge the conceptual gap between Bayesian and frequentist perspectives, and provides a coherent probabilistic framework that may serve as a guiding principle for future work involving uncertainty modeling and data combination strategies.
Before closing, we note possible directions for future work. The hierarchical Bayesian model with hyperpriors on variance parameters [9] has been proposed as a way to address issues like “sub-sampling” [10], where iterative rescaling can bias the combined central value. It would be worthwhile to examine whether this issue can also be resolved within a frequentist approach, and whether the correspondence identified in this Letter continues to hold in that context. In parallel, it would also be valuable to explore how the frameworks discussed here perform when applied to real-world data sets [11], such as those underlying the $W$ boson mass [12], the muon g2g-2 anomaly [4],111111 Applications to the muon g2g-2 have already been considered in Ref. [5]; see also recent developments on the data-theory comparison in Refs. [13, 14]. the neutron lifetime [15] (and references therein), and the $\sigma_{8}$ tension in cosmology [16]. Such applications may offer not only practical validation but also a clearer understanding of the statistical structure behind persistent discrepancies in precision measurements.